# MonoDream: Monocular Vision-Language Navigation with Panoramic Dreaming

### v1

@@ Abstract
1 Introduction
Vision-Language Navigation tasks (Wu et al. 2024; Anderson et al. 2018; Gu et al. 2022) require embodied agents to follow language instructions and navigate to specified targets in 3D environments. Early successful approaches often rely on global perceptual inputs, such as panoramic RGB-D images (Hong et al. 2022; Wang et al. 2023; An et al. 2022), which provide a wide field of view and explicit visual and geometric information. These inputs allow agents to build a more complete understanding of the environment and achieve high navigation success rates.
However, panoramic cameras and depth sensors introduce higher cost, power consumption, added weight and hardware integration complexity, making them less practical in many real-world deployments. Recent research has therefore focused on more lightweight settings where the agent is equipped with only a single forward-facing RGB camera (Zhang et al. 2024b, a; Cheng et al. 2024).
@@ 1 Introduction
To tackle this issue, we propose MonoDream, a lightweight VLA framework that equips monocular VLN agents with a latent imagination capability. Our approach is built upon a key insight: navigation-relevant information, including the predicted actions and the understanding of the global scene, can be encoded into a shared representation space and could be inferred by the agent. This insight is inspired by neuroscientific findings that the human brain reasons the current panoramic scene from partial views (Robertson et al. 2016) and internally simulates upcoming scenes based on intention (Seeber et al. 2025). MonoDream enables monocular agents to learn to infer and complete a holistic understanding of the current and future environments from limited egocentric observations.
@@
2 Related Work
2.1 Vision Language Navigation
Vision-Language Navigation tasks (Nguyen et al. 2019; Wang et al. 2022; Wu et al. 2020; Schumann et al. 2024) challenge an embodied agent to follow natural language instructions and reach a target location within a previously unseen environment. Although early approaches largely focus on discrete navigation settings (Hong et al. 2021; Chen et al. 2021; Liu et al. 2023), where the agent moves along a predefined graph and typically has access to panoramic RGB-D observations, more recent studies (Hong et al. 2022; Wang et al. 2023; An et al. 2022; Dai et al. 2024) have shifted toward more realistic, continuous environments to predict low-level actions, and some works also begin to investigate monocular RGB-D settings.
Recent works based on large models (Liu et al. ; Zhang et al. 2024a; Cheng et al. 2024; Wei et al. 2025) leverage large-scale, RGB-only video models to build monocular VLN systems with enhanced generalization and real-world applicability. However, agents in these works rely on a single forward-facing RGB camera, which poses additional challenges due to the limited field of view and the absence of explicit depth information. Some works (Wang et al. 2024b; Wang, Lee, and Lee 2025) attempt to use neural rendering techniques to recover observations from novel viewpoints from the reconstructed map. However, these methods rely on additional localization and mapping modules, and their performance still falls significantly short compared to approaches that directly utilize the panoramic input.
@@ 2.1 Vision Language Navigation
2.2 Vision Imagination in Embodied Agents
Learning to imagine vision information has become an increasingly popular strategy for improving policy learning in embodied agents, particularly when the agent operates under egocentric and limited monocular observations. In such settings, inferring spatial layouts, geometric structures, and future dynamics beyond the current observation becomes essential for decision-making. Early approaches (Wu et al. 2023; Wang et al. 2024a) often adopt a two-stage pipeline, where future states are predicted using off-the-shelf world models (Zheng et al. 2024; Bar et al. 2025) or neural rendering (Mildenhall et al. 2021; Kerbl et al. 2023) from reconstructed maps. Although simple and modular, these methods are inherently constrained by the accuracy and generalizability of the underlying world models. More recent works (Cen et al. 2025; Tian et al. 2024; Chen et al. 2024) advocate for end-to-end training paradigms that integrate forecasting and action planning within a unified framework. These methods allow the agent to jointly predict future observations and actions, leading to improved performance and generalization. Notably, DreamVLA (Zhang et al. 2025) introduces a future-aware policy that forecasts dynamics, depth, and semantics simultaneously from monocular RGB input, demonstrating the benefits of future modeling in visual language tasks.
@@ 3 Method
3.1 Unified Navigation Representation
@@
The agent’s visual input contains the current view $o_{t}$ and a sampled historical images. To maintain computational efficient while preserving essential historical context, we uniformly sample $N$ frames, denoted as $\mathcal{O}_{t}=\{o_{p_{0}},\dots,o_{p_{N-1}}\}$, from the full sequence of past observations. Each of these images is independently processed by the vision encoder, $\Phi_{\text{vis}}(\cdot)$, which encodes each image into $d$-dimensional feature and these individual image features are collected to form the visual input sequence $E_{\text{vis}}$. $E_{\text{vis}}$ represents the agent’s complete visual context at step $t$ from monocular image sequence.
@@
We define hidden feature $h_{t}$ as the Unified Navigation Representation. To ensure that $h_{t}$ encodes all navigation-relevant information, even under limited monocular observations, we design a multitask training framework. In this framework, MonoDream learns $h_{t}$ to capture both vision-based signals (global layout, geometric structure, and future dynamics) and language-based signals (actions and instructions). Specifically, we jointly train the model with multiple objectives: Latent Panoramic Dreaming (Section 3.2), action prediction and instruction reasoning (Section 3.3).
3.2 Latent Panoramic Dreaming
@@
where =$\mathcal{M}=\{\text{PI},\text{PD}\}$.
@@
3.3 Multi-Task Co-Training
@@
At each time step $t$, the model predicts navigation actions in the natural language form. Specially, the model predicts a sequence of the next three actions $(a_{t},a_{t+1},a_{t+2})$ based on the embedding $h_{t}$, generated from the instruction $I$, current observation $o_{t}$, and navigation history $\mathcal{O}_{t}$ for the sample at step $t$. This setup encourages short-term action forecasting while retaining reactivity to new observations. The training objective for each timestamp $t$ is defined as:
@@
For each training trajectory, we uniformly sample $N$ visual observations to construct the input sequence $O_{t}$. These images are processed by the MonoDream backbone to produce the UNR $h_{t}$. A text decoder is then applied to generate the distribution over instruction tokens based on $h_{T}$, and the loss is computed accordingly:
@@ 4 Implementation Details
All the training data used in our work are collected from simulated environments, including the training splits of R2R-CE (Krantz et al. 2020) and RxR-CE (Ku et al. 2020), as well as additional data collected using the DAgger (Ross, Gordon, and Bagnell 2011) strategy. We first construct step-wise navigation data based on the action annotations provided in R2R-CE and RxR-CE, resulting in 320K and 600K samples respectively. In addition, we construct auxiliary supervision data based on R2R-CE by applying the aforementioned image preprocessing and instruction reasoning tasks. Moreover, following the DAgger strategy (Ross, Gordon, and Bagnell 2011), we further collect 500K step-wise samples from non-oracle trajectories generated in the R2R-CE training environments.
For the training panoramic images in LPD tasks, we adapt the cubemap (Trindade and Raposo 2011) format and split into four canonical directions (left, front, right, and back). For the depth image, we apply logarithmic scaling to the raw depth values to compress large-scale variations, followed by a colormap-based RGB rendering (Itseez 2015). See our supplementary materials for more details of data pre-processing.
@@
We adopt NVILA-lite-2B (Liu et al. 2024) as our base model, which includes a SigLIP vision encoder (Tschannen et al. 2025), a projection module, and a Qwen2-based language model (Bai et al. 2025). Starting from NVILA-lite-2B’s pre-trained weights, we perform supervised fine-tuning (SFT) for our tasks. All components of the model are trainable during fine-tuning. Training is conducted on 8 NVIDIA H20 GPUs for 5 epochs with the learning rate of 1e-5, the warm-up ratio of 0.03, and the batch size of 80. During training and inference, we set the number of future actions to predict ($K$) to 3, and the number of historical frames to sample ($N$) to 8.
Table 1: Comparison of different methods on the R2R Val-Unseen split. Observations used include single RGB camera (S.RGB), depth sensor (Depth) and panoramic view (Pano.). $\dagger$ indicates methods without using LLMs. External data refers to sources beyond the navigation simulator, such as real-world web data, general VQA datasets, and other similar resources.
Method Observation R2R Val-Unseen Training
BEVBert(An et al. 2022) $\checkmark$ $\checkmark$ 4.57 67.0 59.0 50.0 -
ETPNav(An et al. 2024) $\checkmark$ $\checkmark$ 4.71 65.0 57.0 49.0 -
ENP–ETPNav(Liu, Wang, and Yang 2024) $\checkmark$ $\checkmark$ 4.69 65 58 50 -
Seq2Seq(Krantz et al. 2020) $\checkmark$ $\checkmark$ 7.77 37.0 25.0 22.0 -
CMA(Krantz et al. 2020) $\checkmark$ $\checkmark$ 7.37 40.0 32.0 30.0 -
LAW(Raychaudhuri et al. 2021) $\checkmark$ $\checkmark$ 6.83 44.0 35.0 31.0 -
CM2(Georgakis et al. 2022) $\checkmark$ $\checkmark$ 7.02 41.0 34.0 27.0 -
WS-MGMap(Chen et al. 2022) $\checkmark$ $\checkmark$ 6.28 47.0 38.0 34.0 -
sim2real(Wang et al. 2024b) $\checkmark$ $\checkmark$ 5.95 55.8 44.9 30.4 -
NavMorph(Yao, Gao, and Xu 2025) $\checkmark$ $\checkmark$ 5.75 56.9 47.9 33.2 -
NaVid-4D (Liu et al. ) $\checkmark$ $\checkmark$ 5.99 55.7 43.8 37.1 1500K
NaVid (Zhang et al. 2024b) $\checkmark$ 5.47 49.1 37.4 35.9 0K
Uni-NaVid(Zhang et al. 2024a) $\checkmark$ 5.58 53.3 47.0 42.7 2300K
NaVILA(Cheng et al. 2024) $\checkmark$ 5.22 62.5 54.0 49.0 2215K
MonoDream $\checkmark$ 5.45 61.5 55.8 49.1 0K
5 Experimental Results
5.1 Experiment Setup
@@
We evaluate our method on the VLN-CE benchmarks R2R-CE (Krantz et al. 2020) and RxR-CE (Ku et al. 2020) following the standard VLN-CE settings. All the methods are evaluated on the R2R val-unseen split and RxR val-unseen split. The quantitative results are in Figure 2.
@@
We follow the standard VLN evaluation protocol (Krantz et al. 2020; Ku et al. 2020) to evaluate navigation performance for all methods, including success rate (SR), oracle success rate (OSR), success weighted by path length (SPL), and the navigation error from the goal (NE). Among them, SR and SPL are widely regarded as the primary metrics, reflecting the task completion and path efficiency respectively.
5.2 Comparison on VLN-CE Benchmarks
@@
To further evaluate MonoDream’s robustness and long-term ability, we conduct experiments on the RxR-CE Val-Unseen split (Ku et al. 2020), as in Table 4. Compared to R2R-CE, RxR-CE presents significantly more complex challenges due to its longer trajectories and more natural, diverse language instructions. Despite these challenges and the monocular setting, MonoDream achieves state-of-the-art results on the primary metrics (SR and SPL), outperforming strong baselines like Uni-NaVid (Zhang et al. 2024a) and NaVILA (Cheng et al. 2024) while using substantially less training data. These findings demonstrate the advantage of leveraging visual imagination as auxiliary supervision.
@@
We assess the generalization capability of MonoDream on the RxR-CE Val-Unseen split (Table 2). Notably, all the methods are trained without any RxR-CE training data and our method achieves the state-of-the-art performance. This demonstrates the strong transferability of our framework.
@@
Seq2Seq(Krantz et al. 2020) 11.8 5.02 3.51 3.43
CMA(Krantz et al. 2020) 11.7 10.7 4.41 2.47
LAW(Raychaudhuri et al. 2021) 10.87 21.0 8.0 8.0
CM2(Georgakis et al. 2022) 8.98 25.3 14.4 9.2
WS-MGMap(Chen et al. 2022) 9.83 29.8 15.0 12.1
A2NAV(Chen et al. 2023) - - 16.8 6.3
NaVid(Zhang et al. 2024b) 8.41 34.5 23.8 21.2
@@
5.3 Ablation Study
In the ablation study, we separately evaluate the effectiveness of the two auxiliary tasks, Latent Panoramic Dreaming and Instruction Reasoning, as well as the individual contributions of the four sub-tasks in LPD. To ensure efficiency, all ablation experiments are conducted using only the R2R-CE training split and trained for one epoch.
@@ 5.3 Ablation Study
The most substantial improvement comes from LPD. Introducing LPD achieves the best performance across all metrics. These results confirm the complementary benefits of language- and vision-oriented supervision: IR improves semantic goal alignment, while LPD enriches UNR with global, geometric, and predictive spatial cues that are otherwise missing in monocular observations. The synergy between the two auxiliary tasks significantly strengthens the agent’s ability to navigate under limited sensory input.
@@
Table 4: Comparison of different methods on the RxR Val-Unseen split. Observations used include single RGB camera (S.RGB), depth sensor (Depth) and panoramic view (Pano.). $\dagger$ indicates methods without using LLMs.
Method Observation RxR Val-Unseen
S.RGB Depth Pano. NE $\downarrow$ OSR $\uparrow$ SR $\uparrow$ SPL $\uparrow$
BEVBert(An et al. 2022) 4.00 - 68.5 -
ETPNav(An et al. 2024) 5.64 - 54.7 44.8
ENP–ETPNav(Liu, Wang, and Yang 2024) $\checkmark$ 5.51 - 55.27 45.11
Seq2Seq(Krantz et al. 2020) $\checkmark$ $\checkmark$ 11.8 - 13.9 11.9
LAW(Raychaudhuri et al. 2021) $\checkmark$ $\checkmark$ 10.87 21.0 8.0 8.0
CM2(Georgakis et al. 2022) $\checkmark$ $\checkmark$ 12.29 25.3 14.4 9.2
sim2real(Wang et al. 2024b) $\checkmark$ $\checkmark$ 8.79 36.7 25.5 18.1
Uni-NaVid(Zhang et al. 2024a) $\checkmark$ 6.24 55.5 48.7 40.9
NaVILA(Cheng et al. 2024) $\checkmark$ 6.77 - 49.3 44.0
MonoDream (Ours) $\checkmark$ 6.38 55.8 49.4 40.9
@@
Table 5 presents the ablation study of the four auxiliary tasks in LPD, including the latent feature of panoramic RGB image (PI), depth image (DI), future panoramic RGB image (FPI), and future depth image (FDI). We add each component on top of the baseline and evaluate their individual and cumulative contributions to navigation performance.
@@
5.4 Model Efficiency
@@
As shown in Table 6, our method achieves much smaller model size and the fastest inference speed among all compared approaches. Despite introducing auxiliary latent imagination tasks during training, MonoDream is inference-efficient as the fact that auxiliary modules are disabled at test time. This makes MonoDream a practical solution for real-time embodied navigation applications.
@@
Method Params. Time / Step
NaVid (Zhang et al. 2024b) 7B 1.2s
NaVILA (Cheng et al. 2024) 8B 1.2s
MonoDream (Ours) 2B 0.8s
6 Conclusion

### v3

@@ Abstract
Introduction
Vision-Language Navigation tasks (wu2024vision; anderson2018vision; gu2022vision) require embodied agents to follow language instructions and navigate to specified targets in 3D environments. Early successful approaches often rely on global perceptual inputs, such as panoramic RGB-D images (hong2022bridging; wang2023gridmm; an2022bevbert), which provide a wide field of view and explicit visual and geometric information. These inputs allow agents to build a more complete understanding of the environment and achieve high navigation success rates.
However, panoramic cameras and depth sensors introduce higher cost, power consumption, added weight and hardware integration complexity, making them less practical in many real-world deployments. Recent research has therefore focused on more lightweight settings where the agent is equipped with only a single forward-facing RGB camera (zhang2024navid; zhang2024uni; cheng2024navila).
@@ Introduction
To tackle this issue, we propose MonoDream, a lightweight VLA framework that equips monocular VLN agents with a latent imagination capability. Our approach is built upon a key insight: navigation-relevant information, including the predicted actions and the understanding of the global scene, can be encoded into a shared representation space and could be inferred by the agent. This insight is inspired by neuroscientific findings that the human brain reasons the current panoramic scene from partial views (robertson2016neural) and internally simulates upcoming scenes based on intention (seeber2025human). MonoDream enables monocular agents to learn to infer and complete a holistic understanding of the current and future environments from limited egocentric observations.
@@
Related Work
Vision Language Navigation
Vision-Language Navigation tasks (nguyen2019vision; wang2022towards; wu2020towards; schumann2024velma) challenge an embodied agent to follow natural language instructions and reach a target location within a previously unseen environment. Although early approaches largely focus on discrete navigation settings (hong2021vln; chen2021history; liu2023bird), where the agent moves along a predefined graph and typically has access to panoramic RGB-D observations, more recent studies (hong2022bridging; wang2023gridmm; an2022bevbert; dai2024unitedvln) have shifted toward more realistic, continuous environments to predict low-level actions, and some works also begin to investigate monocular RGB-D settings.
Recent works based on large models (liunavid; zhang2024uni; cheng2024navila; wei2025streamvln) leverage large-scale, RGB-only video models to build monocular VLN systems with enhanced generalization and real-world applicability. However, agents in these works rely on a single forward-facing RGB camera, which poses additional challenges due to the limited field of view and the absence of explicit depth information. Some works (wang2024sim; wang2025dynam3d) attempt to use neural rendering techniques to recover observations from novel viewpoints from the reconstructed map. However, these methods rely on additional localization and mapping modules, and their performance still falls significantly short compared to approaches that directly utilize the panoramic input.
@@
Vision Imagination in Embodied Agents
Learning to imagine vision information has become an increasingly popular strategy for improving policy learning in embodied agents, particularly when the agent operates under egocentric and limited monocular observations. In such settings, inferring spatial layouts, geometric structures, and future dynamics beyond the current observation becomes essential for decision-making. Early approaches (wu2023unleashing; wang2024reinforcement) often adopt a two-stage pipeline, where future states are predicted using off-the-shelf world models (zheng2024open; bar2025navigation) or neural rendering (mildenhall2021nerf; kerbl20233d) from reconstructed maps. Although simple and modular, these methods are inherently constrained by the accuracy and generalizability of the underlying world models. More recent works (cen2025worldvla; tian2024predictive; chen2024moto) advocate for end-to-end training paradigms that integrate forecasting and action planning within a unified framework. These methods allow the agent to jointly predict future observations and actions, leading to improved performance and generalization. Notably, DreamVLA (zhang2025dreamvla) introduces a future-aware policy that forecasts dynamics, depth, and semantics simultaneously from monocular RGB input, demonstrating the benefits of future modeling in visual language tasks.
@@
Unified Navigation Representation
@@
The agent’s visual input contains the current view $o_{t}$ and a sampled historical images. To maintain computational efficient while preserving essential historical context, we uniformly sample NN frames, denoted as $\mathcal{O}_{t}=\{o_{p_{0}},\dots,o_{p_{N-1}}\}$, from the full sequence of past observations. Each of these images is independently processed by the vision encoder, $\Phi_{\text{vis}}(\cdot)$, which encodes each image into $d$-dimensional feature and these individual image features are collected to form the visual input sequence $E_{\text{vis}}$. $E_{\text{vis}}$ represents the agent’s complete visual context at step $t$ from monocular image sequence.
@@
We define hidden feature $h_{t}$ as the Unified Navigation Representation. To ensure that $h_{t}$ encodes all navigation-relevant information, even under limited monocular observations, we design a multitask training framework. In this framework, MonoDream learns $h_{t}$ to capture both vision-based signals (global layout, geometric structure, and future dynamics) and language-based signals (actions and instructions). Specifically, we jointly train the model with multiple objectives: Latent Panoramic Dreaming (Section Latent Panoramic Dreaming), action prediction and instruction reasoning (Section Multi-Task Co-Training).
Latent Panoramic Dreaming
@@
where $\mathcal{M}=\{\text{PI},\text{PD}\}$.
@@
Multi-Task Co-Training
@@
At each time step $t$, the model predicts navigation actions in the natural language form. Specially, the model predicts a sequence of the next three actions $(a_{t},a_{t+1},a_{t+2})$ based on the embedding $h_{t}$, generated from the instruction II, current observation $o_{t}$, and navigation history $\mathcal{O}_{t}$ for the sample at step $t$. This setup encourages short-term action forecasting while retaining reactivity to new observations. The training objective for each timestamp $t$ is defined as:
@@
For each training trajectory, we uniformly sample NN visual observations to construct the input sequence $O_{t}$. These images are processed by the MonoDream backbone to produce the UNR $h_{t}$. A text decoder is then applied to generate the distribution over instruction tokens based on $h_{T}$, and the loss is computed accordingly:
@@
Method Observation R2R Val-Unseen Training
BEVBert(an2022bevbert) $\checkmark$ $\checkmark$ 4.57 67.0 59.0 50.0 -
ETPNav(an2024etpnav) $\checkmark$ $\checkmark$ 4.71 65.0 57.0 49.0 -
ENP–ETPNav(liu2024vision) $\checkmark$ $\checkmark$ 4.69 65 58 50 -
Seq2Seq(krantz2020beyond) $\checkmark$ $\checkmark$ 7.77 37.0 25.0 22.0 -
CMA(krantz2020beyond) $\checkmark$ $\checkmark$ 7.37 40.0 32.0 30.0 -
LAW(ray2021language) $\checkmark$ $\checkmark$ 6.83 44.0 35.0 31.0 -
CM2(georgakis2022cross) $\checkmark$ $\checkmark$ 7.02 41.0 34.0 27.0 -
WS-MGMap(chen2022weakly) $\checkmark$ $\checkmark$ 6.28 47.0 38.0 34.0 -
sim2real(wang2024sim) $\checkmark$ $\checkmark$ 5.95 55.8 44.9 30.4 -
NavMorph(yao2025navmorph) $\checkmark$ $\checkmark$ 5.75 56.9 47.9 33.2 -
NaVid-4D (liunavid) $\checkmark$ $\checkmark$ 5.99 55.7 43.8 37.1 1500K
NaVid (zhang2024navid) $\checkmark$ 5.47 49.1 37.4 35.9 0K
Uni-NaVid(zhang2024uni) $\checkmark$ 5.58 53.3 47.0 42.7 2300K
NaVILA(cheng2024navila) $\checkmark$ 5.22 62.5 54.0 49.0 2215K
Aux-Think (wang2025think) $\checkmark$ 5.49 62.9 55.7 48.7 1600K
MonoDream (Ours) $\checkmark$ 5.45 61.5 55.8 49.1 0K
Table 1: Comparison of different methods on the R2R-CE Val-Unseen split. Observations used include single RGB camera (S.RGB), depth sensor (Depth) and panoramic view (Pano.). $\dagger$ indicates methods without using LLMs. External data refers to sources beyond the navigation simulator, such as real-world web data, general VQA datasets, and other similar resources.
@@
All the training data used in our work are collected from simulated environments, including the training splits of R2R-CE (krantz2020beyond) and RxR-CE (ku2020room), as well as additional data collected using the DAgger (ross2011reduction) strategy. We first construct step-wise navigation data based on the action annotations provided in R2R-CE and RxR-CE, resulting in 320K and 600K samples respectively. In addition, we construct auxiliary supervision data based on R2R-CE by applying the aforementioned image preprocessing and instruction reasoning tasks. Moreover, following the DAgger strategy (ross2011reduction), we further collect 500K step-wise samples from non-oracle trajectories generated in the R2R-CE training environments.
For the training panoramic images in LPD tasks, we adapt the cubemap (trindade2011improving) format and split into four canonical directions (left, front, right, and back). For the depth image, we apply log scaling to the raw values to compress large-scale variations, followed by a colormap-based RGB rendering (opencv_library).
@@
We adopt NVILA-lite-2B (liu2024nvila) as our base model, which includes a SigLIP vision encoder (tschannen2025siglip), a projection module, and a Qwen2-based language model (bai2025qwen2). Starting from NVILA-lite-2B’s pre-trained weights, we perform supervised fine-tuning for our tasks. All components of the model are trainable during fine-tuning. Training is conducted on 8 NVIDIA H20 GPUs for 5 epochs with the learning rate of 1e-5 , the warm-up as 0.03, and the batch size of 80. During training and inference, we set the number of future actions to predict ($K$) to 3, and the number of historical frames to sample (NN) to 8.
Experimental Results
Experiment Setup
@@
We evaluate our method on the VLN-CE benchmarks R2R-CE (krantz2020beyond) and RxR-CE (ku2020room) following the standard VLN-CE settings. All the methods are evaluated on the R2R val-unseen split and RxR val-unseen split. The quantitative results are in Figure 2.
@@
We follow the standard VLN evaluation protocol (krantz2020beyond; ku2020room) to evaluate navigation performance for all methods, including success rate (SR), oracle success rate (OSR), success weighted by path length (SPL), and the navigation error from the goal (NE). Among them, SR and SPL are widely regarded as the primary metrics, reflecting the task completion and path efficiency respectively.
Comparison on VLN-CE Benchmarks
@@
To further evaluate MonoDream’s robustness and long-term ability, we conduct experiments on the RxR-CE Val-Unseen split (ku2020room), as in Table 2. Compared to R2R-CE, RxR-CE presents significantly more complex challenges due to its longer trajectories and more natural, diverse language instructions. Despite these challenges and the monocular setting, MonoDream achieves state-of-the-art results on the primary metrics (SR and SPL), outperforming strong baselines like Uni-NaVid (zhang2024uni) and NaVILA (cheng2024navila) while using substantially less training data. These findings demonstrate the advantage of leveraging visual imagination as auxiliary supervision.
@@
We assess the generalization capability of MonoDream on the RxR-CE Val-Unseen split (Table 3). Notably, all the methods are trained without any RxR-CE training data and our method achieves the state-of-the-art performance.
@@
CMA (hong2022bridging) 8.76 - 26.5 22.1
VLNBERT (hong2022bridging) 8.98 - 27.0 22.6
Seq2Seq(krantz2020beyond) 11.8 - 13.9 11.9
sim2real(wang2024sim) 8.79 36.7 25.5 18.1
Uni-NaVid(zhang2024uni) 6.24 55.5 48.7 40.9
NaVILA(cheng2024navila) 6.77 - 49.3 44.0
MonoDream (Ours) 6.38 55.8 49.4 40.9
Table 2: Comparison of different methods on the RxR Val-Unseen split. $\dagger$ indicates methods without using LLMs.
Method RxR Val-Unseen
NE $\downarrow$ OSR $\uparrow$ SR $\uparrow$ SPL $\uparrow$
Seq2Seq(krantz2020beyond) 11.8 5.02 3.51 3.43
CMA(krantz2020beyond) 11.7 10.7 4.41 2.47
LAW(ray2021language) 10.87 21.0 8.0 8.0
CM2(georgakis2022cross) 8.98 25.3 14.4 9.2
WS-MGMap(chen2022weakly) 9.83 29.8 15.0 12.1
A2NAV(chen20232) - - 16.8 6.3
NaVid(zhang2024navid) 8.41 34.5 23.8 21.2
@@
Ablation Study
In the ablation study, we separately evaluate the effectiveness of the two auxiliary tasks, LPD and Instruction Reasoning, as well as the individual contributions of the four sub-tasks in LPD. To ensure efficiency, all ablation experiments are conducted using only the R2R-CE training split and test on R2R-CE val-unseen split only.
@@
The most substantial improvement comes from LPD. Introducing LPD achieves the best performance across all metrics. These results confirm the complementary benefits of language- and vision-oriented supervision: IR improves semantic goal alignment, while LPD enriches UNR with global, geometric, and predictive spatial cues that are otherwise missing in monocular observations.
@@
Table 5 presents the ablation study of the four auxiliary tasks in LPD, including the latent feature of panoramic RGB image (PI), panoramic depth (PD), future panoramic RGB image (FPI), and future panoramic depth (FPD). We add each component on top of the baseline and evaluate their individual and cumulative contributions to navigation performance.
@@
Impact of Prediction Steps in Latent Panoramic Dreaming
We further investigate the effect of predicting the future panoramic feature of different future steps in LPD, and we take FPI as a representative example. As shown in Table 6, predicting a single future step delivers the best performance. Extending the prediction horizon to 2 or 3 future steps leads to performance degradation, due to compounded uncertainty in long-range forecasting under partial observations.
[…]
1 7.03 44.5 39.1 34.2
2 7.89 43.9 34.3 29.1
3 8.77 40.0 29.2 26.2
Table 6: Ablation on the prediction steps in LPD.
Model Efficiency
@@
As in Table 7, our method achieves smaller model size and fastest inference speed among compared approaches. Despite introducing auxiliary latent imagination tasks during training, MonoDream is inference-efficient as auxiliary modules are disabled at test time. This makes MonoDream a practical solution for real-time embodied navigation applications.
Method Params. Time / Step
NaVILA (cheng2024navila) 8B 1.2s
Aux-Think (wang2025think) 8B 1.2s
MonoDream (Ours) 2B 0.8s
@@
Conclusion